OpenStudy (anonymous):

FAN AND MEDAL Which best explains whether a triangle with side lengths 2 in., 5 in., and 4 in. is an acute triangle? A.The triangle is acute because 22 + 52 > 42. B.The triangle is acute because 2 + 4 > 5. C.The triangle is not acute because 22 + 42 < 52. D.The triangle is not acute because 22 < 42 + 52.

OpenStudy (welshfella):

Hint:- in an acute angled triangle c^2 < a^2 + b^2 where c is the longest side

OpenStudy (anonymous):

Is it D?

OpenStudy (welshfella):

No 2 is the shortest side You are looking for the longest side to be isolated

OpenStudy (anonymous):

B.

OpenStudy (welshfella):

which side is the longest?

OpenStudy (anonymous):

5

OpenStudy (welshfella):

right - and each side has to be squared

OpenStudy (welshfella):

acute triangle:- c^2 < a^2 + b^2 obtuse triangle c^2 > a^2 + b^2 right triangle:- c^2 = a^2 + b^2

OpenStudy (welshfella):

B is not correct because the values are not squared. ALSO ITS BETTER TO WRITE SQUARED NUMBERS as 2^2 , 4^2 and 5^2 to avoid confusion 22 is usually take to be 'twenty two'.

OpenStudy (welshfella):

the options are written as the reverse of the equations i gave you so i'll rewrite them in this way acute triangle a^2 + b^2 > c^2 obtuse a^2 + b^2 < c^2 That should help yo choose the right one They are the same as the other ones but are written in reverse

OpenStudy (welshfella):

let a = 2 and = 4 and c = 5. (c is always the longest side).

OpenStudy (welshfella):

* b = 4